# Explicit Motion Handling and Interactive Prompting for Video Camouflaged Object Detection

### v1

@@ 4.2 Comparisons with State-of-the-arts
Table 2: Ablation results on camouflage feeder (CF) and motion collector (MC) modules of EMIP on MoCA-Mask dataset. The best results are highlighted in bold.
@@
4.3 Ablation Studies
We conduct thorough ablation studies by removing or substituting components from the complete EMIP.
Camouflage feeder and motion collector. We first validate the role of the two prompt integration modules, i.e., camouflage feeder and motion collector. As shown in Table 2, we start with the baseline model (#1), which directly uses the segmentation stream to extract single-frame information for prediction, i.e., removing the motion modeling stream of EMIP. Configuration #2 means using only motion information to prompt the segmentation stream, meanwhile discarding segmentation-to-motion prompt. Model #3 is the full EMIP. The integration of motion information is beneficial for the detection of camouflaged objects in video sequences. By incorporating the segmentation-to-motion prompt through the camouflage feeder into the motion stream, we enhance the robustness of motion cues, further resulting in improved performance for segmentation.
Full-tuning v.s. Freezing. Existing models usually address motion modeling by fully tuning all the parameters. In contrast, our EMIP adopts a prompt learning strategy to conduct motion modeling by freezing the motion fundamental model. From the comparisons in Table 3, where both settings are preloaded with pre-trained weights, the superiority of freezing the motion model over full-tuning is evident. This implies the prompt learning strategy for the motion modeling stream can better exploit its potential on limited VCOD data.
Table 3: Different learning strategies for the motion stream.
@@ 4.3 Ablation Studies
Table 4: Different prompt destinations on the motion stream.
@@
Prompt destination on the motion stream. The prompt destination should address where to feed the appearance prompt on the motion fundamental model. In this experiment, we utilize $f^{2}_{t}$ as both segmentation-to-motion prompt and motion collector input. Then we use it to prompt different CNN layers, and the evaluation results are presented in Table 4. It can be seen that prompting with the features after CNN Layer3 achieves superior performance. The underlying reason is that distinct-level features need the corresponding matching prompt to enhance level-specific representations.
Prompt source from the segmentation stream. To explore the most suitable appearance prompt to feed to the motion stream, we conduct another experiment with different prompt features, i.e., $f^{2}_{t}$, $f^{3}_{t}$, and $f^{4}_{t}$. The experimental results are presented in Table 5, where it is evident that utilizing appearance features $f^{2}_{t}$ as prompt yields the best performance. One possible explanation for this phenomenon is that slight motion of a moving object is likely to be missed in lower-resolution feature maps, namely $f^{3}_{t}$ and $f^{4}_{t}$, and therefore degrading the efficacy for prompting to the motion stream.
Table 5: Different prompt sources from the segmentation stream.
@@
Prompt destination on the segmentation stream. We leverage motion information to selectively prompt distinct appearance features or simultaneously prompt all features by introducing additional motion collectors. The results, as presented in Table 6, demonstrate that utilizing motion to prompt feature $f^{2}_{t}$ outperforms other features ($f^{3}_{t}$ or $f^{4}_{t}$) or all features collectively. This implies that semantic features at corresponding hierarchical levels between segmentation and motion tasks can mutually enhance the representation of the original features. Yet, the integration of features from disparate levels, which lack alignment, may compromise the fidelity of the original feature representation.
Table 6: Different prompt destinations on the segmentation stream.
@@
Effectiveness of motion self-supervision. Quantitative and qualitative results are reported to validate the effectiveness of different training strategies for the motion stream. Due to the absence of optical flow ground truth, conducting quantitative analyses for the output optical flow is not feasible. Thus, we evaluate its effectiveness via segmentation/prediction performance. From Table 7, the model with a self-supervised loss achieves better performance, which demonstrates that the prompted motion under self-supervision is beneficial to boosting detection performance.
Fig. 5 visualizes comparisons of our model with different settings in terms of optical flow prediction for camouflaged scenarios. The results show that sometimes the original GMFlow cannot perceive camouflaged moving objects (Fig. 5 (c)), whereas ours that incorporates camouflage information with the prompt learning paradigm can better detect the targets (Fig. 5 (a)). Fig. 5 (d) further shows that the lack of self-supervised loss in our EMIP leads to worse flow prediction. In contrast, full-tuning the GMFlow part (Fig. 5 (b), correspond those results in Table 3) is hardly generalized in new camouflaged scenarios.
@@
Table 7: EMIP without/with the self-supervised loss.
@@
Effectiveness of long-term modeling. As aforementioned in Table 1, by introducing the long-term consistency prompt to the short-term model EMIP, we have further enhanced the performance on all five metrics. Additionally, visual comparisons on the same clip in Fig. 6 show that the incorporation of long-term consistency effectively harnesses temporal consistency, thereby reducing errors in short-term prediction.
Figure 6: Visual comparisons on consecutive video frames.
5 Conclusion
@@
We propose EMIP, an innovative framework for VCOD that explicitly handles motion cues through the utilization of a frozen pre-trained optical flow fundamental model. EMIP adopts a novel two-stream architecture, concurrently addressing camouflaged segmentation and optical flow estimation. The core idea of interaction between these two streams is orchestrated through an interactive prompting mechanism. Experimental results show the paradigm of interactive prompting of EMIP can enhance outputs of both streams, further achieving accurate prediction. In addition, we present an extended version of EMIP, incorporating historical features into the prompt to alleviate short-term prediction errors and enhance overall accuracy. Our contributions not only achieve compelling results on two VCOD benchmarks but also provide fresh insights into addressing the challenging VCOD task. We hope that the proposed framework will serve as a catalyst for inspiring further research in this emerging field.

### v2

@@ IV-D1 Camouflage feeder and motion collector
TABLE II: Ablation results on camouflage feeder (CF) and motion collector (MC) modules of EMIP on MoCA-Mask dataset. The best results are highlighted in bold.
@@
Figure 6: Qualitative results of CF and MC. The red box indicates the ground-truth location of the camouflaged object, while the yellow box represents detected noise. With the incorporation of MC, the camouflaged object can be effectively identified. Moreover, integrating CF further suppresses noise.
IV-D2 Full-tuning v.s. Freezing
Existing models usually address motion modeling by fully tuning all the parameters. In contrast, our EMIP adopts a prompt learning strategy to conduct motion modeling by freezing the motion fundamental model. From the comparisons in Table III, where both settings are preloaded with pre-trained weights, the superiority of freezing the motion model over full-tuning is evident. As can be seen from Fig. 7, full-tuning the motion stream disrupts its ability to effectively model motion, introducing additional noise into camouflaged object prediction. Freezing all motion modeling layers allows the model to leverage robust and general features learned from large-scale motion estimation datasets during pre-training. This strategy ensures that the prompting process concentrates on learning camouflage-specific features, thereby enhancing the efficiency and effectiveness of the optimization process of our EMIP. As demonstrated in Table III, the prompt learning strategy within our framework for the motion modeling stream significantly exploits its potential on limited VCOD data, even in the absence of ground truth for optical flow.
TABLE III: Quantitative comparison of full-tuning and freezing the explicit motion modeling stream.
@@
Figure 7: Qualitative results of full-tuning and freezing the motion stream. The yellow boxes represent false positive predictions.
IV-D3 Prompt destination on the motion stream
The prompt destination is crucial as it determines where to introduce the appearance prompt within the motion fundamental model. In this experiment, we selected $f^{2}_{t}$ to serve dual roles: as the segmentation-to-motion prompt and as the input for the motion collector. Subsequently, we tested its effectiveness by prompting various convolutional neural network (CNN) layers within the model. The specific layers chosen for this evaluation ranged from initial to deeper CNN layers, allowing us to observe the impact at different stages. The evaluation results shown in Table IV reveal that prompting with the features immediately after CNN Layer3 yields the best performance among all tested positions. Fig. 8 shows that injecting camouflage priors as prompts after CNN Layer3, but before motion relationship modeling, effectively mitigates the interference of inaccurate motion noise. This observation suggests that the features extracted by CNN Layer3 strike an optimal balance between low-level details and high-level abstractions. The superior performance can be attributed to the necessity for distinct-level features to be complemented by corresponding prompts that enhance level-specific representations. Essentially, the matching prompt at this stage helps in refining and emphasizing the critical features pertinent to motion and appearance, thereby boosting the overall performance of the model. By ensuring that the prompt is aligned with the level-specific characteristics of the features, our model can more effectively capture and utilize the nuanced information present at this stage. This experiment illustrates the importance of precise prompt positioning in the context of motion and appearance integration.
TABLE IV: Quantitative comparison of different prompt destinations on the motion stream.
@@ IV-D3 Prompt destination on the motion stream
Figure 8: Visualization of different prompt destinations on the motion stream. The red box indicates the ground-truth location of the camouflaged object.
IV-D4 Prompt source from the segmentation stream
To thoroughly investigate the most suitable appearance prompt to feed to the motion stream, we conducted an additional set of experiments using various prompt features, i.e., $f^{2}_{t}$, $f^{3}_{t}$, and $f^{4}_{t}$. The experimental results are summarized in Table V and Fig. 9 , clearly demonstrating that employing appearance features $f^{2}_{t}$ as the prompt leads to superior performance. A plausible explanation for this observation is that subtle movements of objects in motion are likely to be overlooked in lower-resolution feature maps, such as $f^{3}_{t}$ and $f^{4}_{t}$. Consequently, this omission diminishes the efficacy of these features when used as prompts in the motion stream, thereby underscoring the importance of higher-resolution appearance features in capturing fine-grained motion details.
TABLE V: Different prompt sources from the segmentation stream.
@@ IV-D4 Prompt source from the segmentation stream
Figure 9: Qualitative results of different prompt sources from the segmentation stream. The red box denotes the ground-truth location of the camouflaged object, while the yellow box highlights the detected noise.
IV-D5 Prompt destination on the segmentation stream
We harness motion information to selectively prompt distinct appearance features or to simultaneously prompt all features by incorporating additional motion collectors into the model. The empirical results, as presented in Table VI, reveal that utilizing motion to specifically prompt feature $f^{2}_{t}$ yields superior performance compared to prompting either $f^{3}_{t}$ or $f^{4}_{t}$ individually or all features collectively. Fig. 10 illustrates that applying a prompt on $f^{2}_{t}$ enables the model to focus on the most relevant object features while effectively suppressing background noise. However, when prompting all features, including $f^{2}_{t}$, mismatched prompt locations can introduce interference with the original features, ultimately degrading detection performance. This finding suggests that semantic features at corresponding hierarchical levels between segmentation and motion tasks can synergistically enhance the representation of the original features. However, integrating features from different hierarchical levels, which lack proper alignment, may compromise the integrity and fidelity of the original feature representation, thereby negatively affecting overall performance.
TABLE VI: Different prompt destinations on the segmentation stream.
@@ IV-D5 Prompt destination on the segmentation stream
Figure 10: Qualitative results of different prompt destinations on the segmentation stream. The yellow boxes represent false positive detections.
IV-D6 Effectiveness of motion self-supervision
Quantitative and qualitative results are reported to validate the effectiveness of different training strategies for the motion stream. Due to the absence of optical flow ground truth, conducting quantitative analyses for the output optical flow is not feasible. Thus, we evaluate its effectiveness via segmentation/prediction performance. As validated in Table VII, the model with a self-supervised loss achieves better performance, which demonstrates that the prompted motion under self-supervision is beneficial to boosting detection performance.
Fig. 11 visualizes comparisons of our model with different settings in terms of optical flow prediction for camouflaged scenarios. The results show that sometimes the original GMFlow cannot perceive camouflaged moving objects (Fig. 11 (c)), whereas ours that incorporates camouflage information with the prompt learning paradigm can better detect the targets (Fig. 11 (a)). Fig. 11 (d) further shows that the lack of self-supervised loss in our EMIP leads to worse flow prediction. In contrast, full-tuning the GMFlow part (Fig. 11 (b), correspond those results in Table III) is hardly generalized in new camouflaged scenarios.
Fig. 12 shows the loss curves of optical flow estimation (left) and camouflaged object segmentation (right), respectively, and one can see that both losses converge during the training process. Notably, the segmentation loss converges rapidly, while the flow loss converges at a slower rate. The straightforward utilization of a well pre-trained optical flow fundamental model results in an initially low flow loss. As the training proceeds, this flow loss enforces stable optimization of the segmentation-to-motion prompt learning, thereby guaranteeing a more effective prompt towards the motion stream.
@@ IV-D6 Effectiveness of motion self-supervision
Figure 12: Training loss of optical flow and segmentation in our proposed model EMIP. Both losses converge after about 80 epochs.
TABLE VII: Quantitative comparison of our EMIP model with and without the self-supervised loss for optical flow.
@@
IV-E Analyses of segmentation-to-motion prompt
We introduce a segmentation-to-motion prompt strategy utilizing a camouflage feeder to enhance motion estimation. This approach refines the features that generate motion information, ensuring more precise responses in camouflaged regions while minimizing the impact of irrelevant information. To better elucidate the inner mechanism and interpretability of segmentation-to-motion prompt, as illustrated in Fig. 13, we visualize several feature maps of $f^{flow}_{t}$, $\hat{f}^{flow}_{t}$, and $f^{2}_{t}$ in one frame. For a more intuitive presentation, we randomly select 10 consecutive feature maps along the channel dimension. As observed, the lower-level appearance features $f^{2}_{t}$ from the segmentation branch exhibits enhanced responses around the camouflaged regions. After being prompted by appearance features, motion features then exhibit an enhanced response towards camouflaged regions. This response aids in further distinguishing the motion features of camouflaged areas from the background pixels, thereby improving the motion estimation for camouflaged objects.
Figure 13: Visualization of the segmentation-to-motion prompt. The left image represents the current frame. The red rectangles highlights prominent regions in feature maps or the current frame. The solid red lines mean the computation processes of the output optical flow using motion features of two adjacent frames ({$f_{t}^{flow}$,$f_{t-1}^{flow}$} or $\hat{f}_{t}^{flow}$,$\hat{f}_{t-1}^{flow}$}).
IV-F Computational Efficiency
To thoroughly assess model-related parameters and efficiency, we perform a comparative analysis in Table VIII against the previous cutting-edge model SLT-Net [10], under the same GPU configuration. Despite having $\sim$18M more parameters compared to SLT-Net, EMIP exhibits decent improvements in both detection accuracy and frames-per-second (FPS). Furthermore, the long-term variant of EMIP, i.e., EMIP, achieves such improvements with remarkably low amount of fine-tuned parameters (8.5M, less than 8% of the total model parameters), reducing the computational overhead during training.
TABLE VIII: Comparison of model parameters and efficiency with previous SLT-Net, in terms of total parameters, finetuned parameters, and inference speed (frames-per-second, FPS). The best scores are highlighted in bold.
Model Total Params Finetuned Params FPS $\mathcal{S}_{\alpha}\uparrow$ $F_{\beta}^{w}\uparrow$ $\mathcal{M}\downarrow$
SLT-Net 82.38M 82.38M 5.5 0.637 0.304 0.027
EMIP 100.86M 96.06M 7.8 0.669 0.374 0.017
EMIP 109.04M 8.50M 6.2 0.675 0.381 0.015
IV-G Generalization Ability on VSOD/VOS
To thoroughly demonstrate the generalizability of EMIP, we conducted extended evaluations using the well-known VSOD/VOS datasets DAVIS16 [70], FBMS [71], ViSal [72] and SegV2 [73]. The results, summarized in Table IX, include a comprehensive quantitative comparison against several recently published state-of-the-art methods for VSOD and VOS tasks. These methods represent a broad spectrum of current advancements in the field. Additionally, we present visual comparisons in Fig. 14 to further illustrate the efficacy of our approach. Our model, EMIP, consistently excels in capturing fine contour details across various scenarios. For instance, it accurately delineates the foot of a dog in the 3rd row, the tail of a horse in the 4th row, and the wheel of a motorcycle in the 5th row. These detailed visual comparisons highlight our model’s ability to handle intricate object boundaries and maintain high fidelity in the segmentation process. The results clearly demonstrate that, although EMIP is specifically designed to tackle video camouflage scenes, it also performs exceptionally well in more general video segmentation tasks. This adaptability underscores the robustness and versatility of our approach, making it suitable for a wide range of applications beyond its original design scope.
[…]
Method DAVIS16 FBMS ViSal SegV2
$\mathcal{S}_{\alpha}\uparrow$ $\mathcal{F}\uparrow$ $\mathcal{M}\downarrow$ $\mathcal{S}_{\alpha}\uparrow$ $\mathcal{F}\uparrow$ $\mathcal{M}\downarrow$ $\mathcal{S}_{\alpha}\uparrow$ $\mathcal{F}\uparrow$ $\mathcal{M}\downarrow$ $\mathcal{S}_{\alpha}\uparrow$ $\mathcal{F}\uparrow$ $\mathcal{M}\downarrow$
SCOM[47]${}_{\text{TIP'2018}}$ 0.832 0.783 0.048 0.794 0.797 0.079 0.762 0.831 0.122 0.815 0.764 0.030
MBNM[48]${}_{\text{ECCV'2018}}$ 0.887 0.861 0.031 0.857 0.816 0.047 0.898 0.883 0.020 0.809 0.716 0.026
PDBM[49]${}_{\text{ECCV'2018}}$ 0.882 0.855 0.028 0.851 0.821 0.064 0.907 0.888 0.032 0.864 0.800 0.024
SRP[50]${}_{\text{TIP'2019}}$ 0.662 0.660 0.070 0.648 0.671 0.134 - 0.752 0.092 - 0.683 0.095
MESO[51]${}_{\text{TMM'2019}}$ 0.718 0.660 0.070 0.635 0.618 0.134 - - - - - -
LTSI[87]${}_{\text{TIP'2019}}$ 0.876 0.850 0.034 0.805 0.799 0.087 0.922 0.909 0.027 0.827 0.862 0.028
RSE[88]${}_{\text{TCSVT'2019}}$ 0.748 0.698 0.063 0.670 0.652 0.128 - - - - - -
SSAV[89]${}_{\text{CVPR'2019}}$ 0.893 0.861 0.028 0.879 0.865 0.040 0.943 0.939 0.020 0.851 0.801 0.023
RCR[52]${}_{\text{ICCV'2019}}$ 0.886 0.848 0.027 0.872 0.859 0.053 - - - - - -
CAS[90]${}_{\text{TNNLS'2020}}$ 0.873 0.860 0.032 0.856 0.863 0.056 - - - 0.820 0.847 0.029
PCSA[91]${}_{\text{AAAI'2020}}$ 0.902 0.880 0.022 0.868 0.837 0.040 0.946 0.940 0.017 0.865 0.810 0.025
DFNet[92]${}_{\text{ECCV'2020}}$ 0.899 0.018 - 0.833 0.054 - 0.927 0.017 - - -
[…]
UFO [93]${}_{\text{TMM'2023}}$ 0.874 0.797 0.032 0.868 0.803 0.041 0.940 0.914 0.012 0.836 0.746 0.057
MAMNet [57]${}_{\text{TIP'2024}}$ 0.897 0.877 0.020 0.894 0.883 0.032 0.947 0.948 0.012 0.886 0.850 0.014
EMIP (Ours) 0.908 0.902 0.016 0.891 0.887 0.032 0.950 0.950 0.012 0.891 0.862 0.013
Figure 14: Visual comparisons of EMIP with four state-of-the-art VOS/VSOD methods. Red rectangles indicate challenging regions on which our EMIP excels.
IV-H Failure Cases
In the 1st scenario depicted in Fig. 15, the camouflaged object is occluded. In this context, the detection outcome encompasses the occlusion object due to the significant resemblance across the camouflaged object, occlusion entity, and background. Another scenario is shown in the 2nd row of the figure. The displayed image represents a frame among initial frames of a sequence, lacking motion information and posing a challenge even for human visual perception, let alone learning models. For the 3rd scenario, the prominent green leaves in the foreground act as strong distractors and the camouflaged object is easily mistaken as part of the foreground foliage. It is important to note that these challenges are not unique to our model. Even state-of-the-art methods from previous research, such as PNS-Net [84] and SLT-Net [10], encounter similar difficulties, as shown in Fig. 15. Consequently, addressing these specific challenges remains a crucial direction for future research in the field of VCOD.
Figure 15: Some failure cases of EMIP and two most recent methods.
V Conclusion
@@
We propose EMIP, an innovative framework for VCOD that explicitly handles motion cues through the utilization of a frozen pre-trained optical flow fundamental model. EMIP adopts a novel two-stream architecture, concurrently addressing camouflaged segmentation and optical flow estimation. The core idea of interaction between these two streams is orchestrated through an interactive prompting mechanism. Experimental results show that the paradigm of interactive prompting of EMIP can enhance the outputs of both streams, further achieving accurate prediction. Comprehensive ablation studies and in-depth discussions validate the key components of EMIP. In addition, we present an extended version of EMIP, incorporating historical features into the prompt to alleviate short-term prediction errors and enhance overall accuracy. Moreover, the proposed EMIP is extended to the general video object segmentation task, consistently delivering improved performance and validating its generalizability and adaptability. Our contributions not only achieve compelling results on two VCOD benchmark datasets, but also provide fresh insights into addressing the challenging VCOD task. We hope that the proposed framework could serve as a catalyst for inspiring further research in this emerging field. We believe that making controllable and adjustable optimization prompts for fundamental models presents an intriguing avenue for future investigation.